OpenStudy (anonymous):

How can 6/36 be reduced to it's simplest form of 1/6??

OpenStudy (anonymous):

divide both the numerator and denominator by 6. or factor out the 6: \[\frac{ 6 }{ 36 } = \frac{ 6(1) }{ 6(6) } = \frac{ 6 }{ 6 } \times \frac{ 1 }{ 6 } = 1 \times \frac{ 1 }{ 6}\]

OpenStudy (anonymous):

ah ok thanks. A previous example showed me to reduce a fraction by subtracting both the Numerator & Denominator to find a number that will go into them both. Then divide the new number by the numerator and denominator to have a new reduced fraction.